# On the local thermodynamic relations in relativistic spin hydrodynamics

### v1

@@ Acknowledgements.
R.S. acknowledges the kind hospitality and support of INFN Firenze, Department of Physics and Astronomy, University of Florence, and ECT* Trento, where part of this work has been carried out. R.S. is also supported partly by a postdoctoral fellowship of West University of Timișoara, Romania. We thank the Galileo Galilei Institute for Theoretical Physics for the hospitality and the INFN for partial support during the completion of this work.
@@ Appendix A Energy-momentum tensor for free massless Dirac field
For the case of global equilibrium with rotation and acceleration, we have the general Belinfante form of energy-momentum tensor for massless fermions as [34, 35]
$T^{\mu\nu}_{\rm B}$ $=$ $\rho u^{\mu}u^{\nu}-p\Delta^{\mu\nu}+Ww^{\mu}w^{\nu}+A\alpha^{\mu}\alpha^{\nu}+G^{l}l^{\mu}l^{\nu}$ (60)
@@
where
$\rho$ $=$ $T^{4}\frac{7\pi^{4}+30\pi^{2}\zeta^{2}+15\zeta^{4}}{60\pi^{2}}-\frac{a^{2}T^{2}(\pi^{2}+3\zeta^{2})}{24\pi^{2}}-\frac{17a^{4}}{960\pi^{2}}$
@@
$p$ $=$ $\frac{\rho}{3},\quad G^{l}=-\frac{2\,T^{4}}{27\pi^{2}},\quad\mathbb{A}=\mathbb{W}=G^{\alpha}=G^{w}=0.$
@@ Appendix B Coefficients of the energy-momentum tensor for free massive Dirac field
The coefficients $\rho$, $U_{\alpha}$, $U_{w}$, $n$, $N_{\alpha}$, $N_{w}$, $A$, $W$, $p$, $D_{\alpha}$, $D_{w}$, $G$, $G_{V}$, and $W^{A}$ of the stress-energy tensor, the vector and axial currents for the massive Dirac fields are given by the general expressions [39]:
$\rho$ $=$ $-\frac{2m^{4}}{\pi^{2}}\Sigma\left[\frac{K_{3}(nx)}{nx}-\frac{K_{2}(nx)}{(nx)^{2}}\right]\cosh{(n\,\zeta)},$
$p$ $=$ $-\frac{2m^{4}}{\pi^{2}}\,\Sigma\left[\frac{K_{2}(nx)}{(nx)^{2}}\right]\cosh(n\,\zeta),$
@@
$n$ $=$ $-\frac{2m^{3}}{\pi^{2}}\Sigma\left[\frac{K_{2}(nx)}{nx}\right]\sinh{(n\,\zeta)},$
@@
where $\Sigma\equiv\Sigma_{n=1}^{\infty}\left(-1\right)^{n+1}$, $\zeta=\mu/T$, and $x=m/T$.

### v3

@@ Acknowledgements.
R.S. acknowledges the kind hospitality and support of INFN Firenze, Department of Physics and Astronomy, University of Florence, and ECT* Trento, where part of this work has been carried out. R.S. is also supported partly by a postdoctoral fellowship of West University of Timișoara, Romania. F.B. is partly supported by the project PRIN2022 Advanced Probes of the Quark Gluon Plasma funded by ”Ministero dell’Università e della Ricerca”, Italy. We thank the Galileo Galilei Institute for Theoretical Physics for the hospitality and the INFN for partial support during the completion of this work. We are grateful to Dirk Rischke for very useful comments.
@@ Appendix A Energy-momentum tensor for free massless Dirac field
For the case of global equilibrium with rotation and acceleration, the exact form of the Belinfante energy-momentum tensor for massless fermions, normalized with Minkowski vacuum subtraction was derived in refs. [34, 35]:
$\tilde{T}^{\mu\nu}_{\rm B}$ $=$ $\bar{\rho}\,u^{\mu}u^{\nu}-\bar{p}\Delta^{\mu\nu}+Ww^{\mu}w^{\nu}+A\alpha^{\mu}\alpha^{\nu}+G^{l}l^{\mu}l^{\nu}$ (67)
@@
where the coefficients read:
$\bar{\rho}$ $=$ $T^{4}\frac{7\pi^{4}+30\pi^{2}\zeta^{2}+15\zeta^{4}}{60\pi^{2}}-\frac{a^{2}T^{2}(\pi^{2}+3\zeta^{2})}{24\pi^{2}}-\frac{17a^{4}}{960\pi^{2}}$
@@
$\bar{p}$ $=$ $\frac{\bar{\rho}}{3},\quad G^{l}=-\frac{2\,T^{4}}{27\pi^{2}},\quad\mathbb{A}=\mathbb{W}=G^{\alpha}=G^{w}=0\,$,
with $\zeta$ being the reduced chemical potential, $\zeta=\mu/T$.
@@ Appendix B Coefficients of the energy-momentum tensor for free massive Dirac field
The coefficients $\bar{\rho}$, $U_{\alpha}$, $U_{w}$, $\bar{n}$, $N_{\alpha}$, $N_{w}$, $A$, $W$, $\bar{p}$, $D_{\alpha}$, $D_{w}$, $G$, $G_{V}$, and $W^{A}$ in the equations (47), (48) and (49) read [39]:
$\bar{\rho}$ $=$ $-\frac{2m^{4}}{\pi^{2}}\Sigma\left[\frac{K_{3}(nx)}{nx}-\frac{K_{2}(nx)}{(nx)^{2}}\right]\cosh{(n\,\zeta)},$
$\bar{p}$ $=$ $-\frac{2m^{4}}{\pi^{2}}\,\Sigma\left[\frac{K_{2}(nx)}{(nx)^{2}}\right]\cosh(n\,\zeta),$
@@
$\bar{n}$ $=$ $-\frac{2m^{3}}{\pi^{2}}\Sigma\left[\frac{K_{2}(nx)}{nx}\right]\sinh{(n\,\zeta)},$
@@
where $\Sigma\equiv\Sigma_{n=1}^{\infty}\left(-1\right)^{n+1}$ and $x=m/T$.